OpenStudy (anonymous):

For the function f(x) = x + 4, what is the ordered pair for the point on the graph when x = 3p? (x, x + 4) (x, 3p + 4) (3p, x + 4) (3p, 3p + 4) Any help? :)

OpenStudy (anonymous):

f(3p) = x + 4 f(3p) = 3p + 4 (3p, 3p+4) x = 3p y= 3p + 4 2

OpenStudy (anonymous):

can you explain?...

OpenStudy (anonymous):

I just showed work

OpenStudy (anonymous):

yes...ooh wait nevermind lol, i didn't get it at first but then i re-read it. Thanks so much! :)

OpenStudy (anonymous):

anytime

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

so you know how to fan people medal people and close questions?

OpenStudy (anonymous):

close questions yes, fan and medal yeaaah no. any help?

OpenStudy (anonymous):

well do you see the best response button next to this comment

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

click that

OpenStudy (anonymous):

ok i did that

OpenStudy (anonymous):

to fan click on my picture and a box pops up. click become a fan. then you can also write a tesimonial

OpenStudy (anonymous):

testimonial or no?

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

:D you know how to do it

OpenStudy (anonymous):

Yeah thanks lol now i don't feel stupid

OpenStudy (anonymous):

haha your welcome

OpenStudy (anonymous):

:)